# Organizing Unstructured Image Collections using Natural Language

### v1

@@ 3.2 Semantic Grouper
The automatically discovered criteria $\mathcal{R}$ serve as thematic indicators for revealing the semantic substructures (or clusters) within the image set $\mathcal{D}$. To uncover these substructures, as shown in Fig. 3(right), the Grouper takes $\mathcal{D}$ and each criterion $R_{l}$ as inputs. It then discovers cluster names $\mathcal{C}_{l}=\{s_{k}^{l}\}_{k=1}^{K_{l}}$ conditioned on $R_{l}$ and assigns each image to its corresponding cluster. The core design of the Grouper focuses on aligning semantic substructure discovery with a specified partitioning criterion. Similar to the Proposer, we also explore three distinct approaches for the Grouper.
@@
Furthermore, as clusters under a given criterion can be formed at varying semantic granularities, depending on user preferences, we have designed our Grouper to cluster $\mathcal{D}$ at three levels of granularity: coarse, middle, and fine-grained. This design enables TeDeSC to provide new insights into the data at different granularities. For example, for the criterion food Cuisine, TeDeSC can organize images at a coarse-grained continental level (e.g., “European” or “Asian”), a middle-grained regional level (e.g., “Mediterranean” or “Southeast Asian”), or a fine-grained national level (e.g., “Italian” or “Thai”).
@@
Image-based Grouper: Given a target criterion $R_{l}\in\mathcal{R}$, we first prompt a LLM to generate a question $q_{l}$ specific to $R_{l}$-i.e., for criterion Mood the generated question is: “What mood is conveyed by this image? Answer with an abstract, common, and specific category name, respectively.” We then use $q_{l}$ to guide a visual question answering (VQA) model to infer semantic cluster names for each image as $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{VQA}(\mathbf{x}_{n},q_{l})$. Consequently, by aggregating the cluster assignment results at each granularity level across $\mathcal{D}$, we can derive multi-granularity semantic substructures.
Tag-based Grouper: Given a criterion, we prompt a LLM to generate a list of common, middle-grained tags specific to that criterion (e.g., “Recreational facility”). Following Liu et al. (2024d), we then obtain coarse- and fine-grained tags by querying the LLM to generate super- and sub-categories (e.g., “Indoor” and “Climbing gym”) for each middle-grained tag. Unlike lexical databases such as WordNet or ConceptNet (Speer et al., 2017), which do not support free-form input and may be limited in accommodating certain criteria, tag synthesis using a LLM provides a flexible and reliable alternative. Subsequently, we use an image tagger to assign the most relevant tag from the candidate tags at each granularity to each image, resulting in multi-granularity substructures after aggregation.
@@
Caption-based Grouper (Main): We prompt a MLLM to generate captions that specifically focus on the target criterion for each image, as $e_{n}^{l}=\text{MLLM}(\mathbf{x}_{n},R_{l})$. Next, we use a LLM in a three-step process to assign images to clusters at multiple granularity levels: i) Initial Naming: First, we prompt the LLM to assign a class name to each caption as $s_{n}^{l}=\text{LLM}(e_{n}^{l},R_{l})$, resulting in an initial set of names $\mathcal{S}_{\text{init}}^{l}$ of size $N$; ii) Multi-granularity Cluster Refinement: Using these initial names as basis, we prompt the LLM to refine them into three structured granularity levels: $\left(\mathcal{S}_{\text{coarse}}^{l},\mathcal{S}_{\text{mid}}^{l},\mathcal{S}_{\text{fine}}^{l}\right)=\text{LLM}(\mathcal{S}_{\text{init}}^{l},R_{l})$. These structured names serve as candidates for cluster assignment; iii) Final Assignment: Each image is then assigned a class name from each granularity level based on its caption, as $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{LLM}(e_{n}^{l},\mathcal{S}_{\text{coarse}},\mathcal{S}_{\text{mid}},\mathcal{S}_{\text{fine}})$. Experiments in Sec. 4.2 show that the Caption-based Grouper performs the best; thus we use it as our main method.
@@ 4 Experiments
Evaluation metrics for Criteria Discovery: Given the discovered criteria set $\mathcal{R}$ of the proposer, we assess its quality from two dimensions: i) Comprehensiveness: We compute True Positive Rate (TPR) (Csurka et al., 2024) for the predicted criteria relative to the annotated ones as $\text{TPR}=\frac{\mid\mathcal{R}\cap\mathcal{Y}\mid}{|\mathcal{Y}|}$, to assess to what extent the predicted set covers the ground-truth set $\mathcal{Y}$; ii) Diversity: We compute the pairwise semantic similarity between the predicted criteria $\mathcal{R}$, using Sentence-BERT (Reimers & Gurevych, 2019), and convert it into a diversity measure by subtracting the similarity from 1. The final score is the average of these pairwise values, reflecting how well the proposed criteria avoid redundancy and capture distinct aspects of the data—i.e., criteria such as Location and Place are nearly identical and provide little additional insight. A higher score indicates better diversity. It is important to note that the number of valid grouping criteria is subjective and potentially unlimited, making False Positives difficult to define. Thus, we use TPR as the primary evaluation metric.
Evaluation metrics for substructure uncovering: Following Conti et al. (2023) and Liu et al. (2024e), for each criterion and its substructure uncovered by the grouper, we evaluate its alignment with the substructure defined by the ground-truth labels along two dimensions: i) Semantic Consistency: For each image, we measure the semantic similarity between its assigned cluster name and the ground-truth label under the current criterion using Sentence-BERT. The average similarity across the dataset is reported as Semantic Accuracy (SAcc), reflecting how well the predicted substructure semantically aligns with the ground-truth. ii) Structural Consistency: We compute the clustering accuracy (CAcc) between the predicted and ground-truth substructures (clusterings) to evaluate how accurately the predicted structure groups images. By combining CAcc and SAcc, we evaluate the overall alignment of the predicted substructure with the ground-truth across various criteria.
Evaluation strategy: Since the granularity of the ground-truth annotations is unknown during evaluation, we choose the substructure from the three predicted granularities that achieves the highest CAcc relative to the ground-truth for assessing our TeDeSC. Unlike methods (Kwon et al., 2024; Yao et al., 2024), which rely on the ground-truth number of clusters to create flat clusterings that perfectly match the ground-truth granularity, we believe our approach offers a fair and practical evaluation strategy for systems capable of providing multi-granularity output.
Figure 4: Comparison of Criteria Proposers. (a) Comprehensiveness: TPR performance of each proposer evaluated on Basic and Hard ground-truth criteria. (b) Diversity: Diversity scores of the proposed criteria for each proposer. (c) Influence of Image Quantity: TPR performance of the caption-based proposer on Hard ground-truth criteria across varying image scales used for criteria discovery. See expanded results in App. E.1.
@@ 4.1 Study of the Criteria Proposer
Expanding ground-truth criteria for comprehensive evaluation: The newly introduced benchmarks provide up to four ground-truth criteria to evaluate the predicted criteria, but for image collections with high scene complexity (e.g.,, COCO-4c), these are still insufficient to cover all valid grouping options. To address this, we expanded the ground-truth criteria for each of the six benchmarks in Sec. 2 by using human annotators to identify additional criteria. This resulted in expanded sets that have: {10, 4, 11, 7, 17, 11} distinct criteria for {Fruit-2c, Card-2c, Action-3c, Clevr-4c, COCO-4c, Food-4c}. We refer to the per-image annotated criteria set (see Fig. 2) as Basic ground truth and the expanded set as Hard. See App. B.2 for the exact criteria and annotation details.
Which Criteria Proposer is best? Our comparative results, presented in Fig. 4(a-b), show that the Caption-based Criteria Proposer consistently outperforms its counterparts by a significant margin across all six benchmarks. As shown in Fig. 4(a), the caption-based proposer achieves an average of 100% TPR on the Basic criteria set and 75.1% TPR on the Hard set across the benchmarks. Its superior performance is particularly evident under the Hard criteria, where it surpasses the second-best tag-based proposer by +32.2% TPR. Meanwhile, the criteria predicted by the caption-based proposer also show greater diversity in Fig. 4(b), coming closest to the diversity of human-annotated criteria.
Although the tag-based proposer performs well on object-centric benchmarks like Clevr-4c or Food-4c, it significantly lags behind the caption-based proposer on more challenging benchmarks such as COCO-4c and Action-3c, which feature more complex, realistic images. This discrepancy arises because image tagging is a more object-centric way of representing image semantics, potentially missing the rich context depicted in images. Conversely, while the image-based proposer outperforms the tag-based proposer on COCO-4c and Action-3c, it is less effective overall because it is limited to reasoning over a small subset of images and loses visual details when combining images into a grid.
Studying the influence of image quantity on Criteria Discovery: Fig. 4(c) shows the TPR performance of the caption-based proposer across varying image scales (100% to only 1 image) used for criteria proposing, tested on the Hard criteria sets of six benchmarks. Interestingly, satisfactory performance is achieved with just a few images in object-centric benchmarks like Card-2c and Clevr-4c. In fact, even a single image often provides sufficient information for reasonable criteria discovery in these object-centric datasets, where a brief glimpse often represents the entirety adequately. For example, seeing one playing card allows the proposer to easily suggest criteria like “Rank” and “Suit” in the Card-2c dataset. However, this does not hold for more complex datasets like COCO-4c, Food-4c, and Action-3c, which contain diverse and realistic scenarios. Here, a reduction in image scale leads to a clear drop in TPR performance, as these datasets require a larger set of images to capture their intricate and varied thematic criteria. Since TeDeSC operates without prior knowledge of the dataset, we default to using the entire dataset to ensure comprehensive criteria discovery.
Figure 5: Comparison of the Semantic Groupers. We show CAcc, SAcc, their Harmonic Mean (HM) evaluated on the basic criteria. The Caption-based Proposer is used by all groupers. See expanded results in App. E.2.
4.2 Study of the Semantic Grouper
Which Semantic Grouper is best? To determine which grouper best uncovers the substructures indicated by the discovered criteria and aligns most closely with the ground truth, we use the Harmonic Mean (HM) of the CAcc and SAcc for each criterion. This allows us to identify the best overall performer for each criterion. In the comparison, to further illustrate the relative performance position of our framework, we first establish a pseudo upper-bound reference by leveraging the zero-shot classification capabilities of CLIP ViT-L14 (Radford et al., 2021) with ground-truth labels for each criterion, where the grouping criteria, cluster names, and number of clusters are all known. Additionally, we apply KMeans (Ahmed et al., 2020) with the ground-truth number of clusters, using robust visual features from CLIP ViT-L14, DINOv1-B/16 (Li et al., 2022), and DINOv2-G/14 (Oquab et al., 2023) to establish baseline positions for clustering performance.
Our comparative analysis, presented in Fig. 5, shows that the Caption-based Semantic Grouper performs best, ranking first in 12 out of 19 tested criteria based on the HM, across six benchmarks. Specifically, in terms of CAcc, the caption-based grouper achieves an average of 59.9% across all tested criteria, closely aligning with the pseudo-upper-bound average of 58.1% and significantly outperforming feature representation-driven baselines, with the highest being DINOv2 at 40.6% CAcc. This demonstrates the effectiveness of our text-driven approach, as the cluster names inferred by the caption-based provide robust guidance for grouping. For SAcc, while the caption-based grouper falls short of the upper-bound reference average of 74.2%, where exact ground-truth class names are used, it still achieves a notable average of 60.5% across all tested criteria, surpassing its counterparts. To some extent, this gap in semantic consistency is expected due to the open nature of the semantic space—i.e., names like “Joyful”, “Happy”, and “Cheerful” are often used to describe the same Mood, yet they do not exhibit full semantic equivalence. The BLIP-2 image-based grouper also shows commendable performance, ranking first in 5 of the 19 tested criteria. This suggests that the criterion-customized prompts generated by LLM help BLIP-2 accurately label relevant visual content. However, its per-image labeling approach may lead to noisy clusters due to a lack of overall alignment between them. The tag-based grouper lags behind in all benchmarks likely due to a mismatch between the generated tags and the concepts in the dataset.
@@ 4.2 Study of the Semantic Grouper
Table 2: Comparison with criterion-conditioned clustering methods. For each benchmark, we report the average CAcc (%) and SAcc (%) across all criteria. We provide CLIP ViT-L/14 zero-shot performance as the pseudo upper-bound reference (UB). See expanded per-criterion results in App. E.3.
@@
MMaP 33.9 - 43.8 - 62.8 - 60.6 - 36.9 - 51.0 - 41.3 -
TeDeSC (Ours) 51.2 48.4 48.1 64.9 64.9 54.3 68.3 60.6 73.3 84.3 65.1 61.1 53.0 53.4
@@
Figure 6: Ablation study of multi-granularity refinement. For each benchmark, we show average CAcc and SAcc across all criteria. See expanded per-criterion numerical results in App. E.4.
Necessity of multi-granularity cluster refinement: To validate the effectiveness of the multi-granularity cluster refinement design, we design control experiments with our Caption-based Grouper, using three different methods for constructing cluster names to organize images: i) Initial Names: using the initially assigned names as the final output, ii) Flat Refinement: prompting the LLM to refine the initial names into a flat list with a unified granularity, and our iii) Multi-granularity Refinement. We then compare the performance of these approaches in Fig. 6. We observe that both refinement methods significantly improve clustering accuracy compared to using noisy initial names, highlighting the importance of having granularity-consistent cluster names for accurately revealing substructures. Additionally, our proposed multi-granularity method surpasses flat refinement, as it enables the grouper to generate clustering results at different granularities, providing greater flexibility in aligning with the label granularity-the user-preferred level of detail.
Additional studies: In App. F, we provide further studies on: i) Fine-grained Scenario: We examine how to integrate our framework with more advanced cross-modal Chain-of-Thought prompting strategies to better handle fine-grained criteria. ii) Sensitivity Analysis: We analyze the system sensitivity to various MLLMs and LLMs. Additionally, Sec. G offers a Qualitative Analysis by visualizing clustering results for different criteria, and Sec. H includes a Failure Case Analysis.

### v4

@@ 4.2 Semantic Grouper
We explore three Grouper designs that partition the dataset $\mathcal{D}$ into semantic clusters by conditioning the partition on each criterion $R_{l}\in\mathcal{R}$. The end result is a partition (or substructure) $\mathcal{O}_{l}=\{\mathcal{C}_{k}^{l}=\bigl{(}s_{k}^{l},\mathcal{D}_{k}^{l}\bigr{)}\}_{k=1}^{K_{l}}$, where an image is not only assigned to a cluster, but also has an associated semantic name (e.g., “Climbing gym” instead of a cluster index), making it human interpretable. Each design differs in how we condition the predictor to incorporate the grouping criterion.
@@
Since data can be clustered at varying granularity, we design the Grouper to cluster images at three granular levels: coarse, middle, and fine. As an example, under the Cuisine criterion, the Grouper can cluster $\mathcal{D}$ at a coarse continental level (“European”, “Asian”), a middle regional level (“Mediterranean”, “Southeast Asian”), or a fine national level (“Italian”, “Thai”).
@@
Image-based Grouper. We task an LLM to generate a question $q_{l}$ conditioned on $R_{l}$. For e.g., for $R_{l}=\leavevmode\nobreak$Mood, the $q_{l}=$ “What mood does this image convey? Return an abstract, common, and specific category.” We then query a visual question answering (VQA) model (Li et al., 2023a) model to output a coarse-, mid-, and fine-grained label for every image as $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{VQA}(\mathbf{x}_{n},q_{l})$. Subsequently, aggregating these labels over $\mathcal{D}$ at each level yields multi-granularity substructures.
Tag-based Grouper: We first prompt a LLM to generate a list of middle-grained tags (e.g., “Recreational facility”) conditioned on a given criterion $R_{l}$. Following Liu et al. (2024d), we then query the LLM to generate super- and sub-categories (e.g., “Indoor” and “Climbing gym”) for each middle-grained tag, obtaining coarse- and fine-grained tags. Given the vocabulary of tags, we use CLIP (Radford et al., 2021) to tag an image with the most likely label, one at each granular level.
@@
Caption-based Grouper (Main): We condition a MLLM with a given criterion $R_{l}$ to generate a set of captions for each image as $e_{n}^{l}=\text{MLLM}(\mathbf{x}_{n},R_{l})$. Next, we use a LLM in a three-step process to group images to clusters at multiple granularity levels: i) Initial Naming: The LLM assigns a class name to each caption as $s_{n}^{l}=\text{LLM}(e_{n}^{l},R_{l})$, producing an initial set of names $\mathcal{S}_{\text{init}}^{l}$; ii) Multi-granularity Cluster Refinement: The LLM refines $\mathcal{S}_{\text{init}}^{l}$ into three structured granularity levels as $\left(\mathcal{S}_{\text{coarse}}^{l},\mathcal{S}_{\text{mid}}^{l},\mathcal{S}_{\text{fine}}^{l}\right)=\text{LLM}(\mathcal{S}_{\text{init}}^{l},R_{l})$, which serve as candidate cluster names; iii) Final Assignment: LLM assigns each image $\mathbf{x}_{n}$ to a cluster by linking its criterion-specific caption to the structured class names each granularity level as $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{LLM}(e_{n}^{l},\mathcal{S}_{\text{coarse}},\mathcal{S}_{\text{mid}},\mathcal{S}_{\text{fine}})$.
@@ 5 Experiments
Evaluation metric for criteria discovery. We assess criterion discovery performance of the proposers using the True Positive Rate (TPR) (Csurka et al., 2024), defined as $\text{TPR}=\frac{\mid\mathcal{R}\cap\mathcal{Y}\mid}{|\mathcal{Y}|}$, i.e., the fraction of ground-truth criteria $\mathcal{Y}$ recovered by the predicted set $\mathcal{R}$. Because the space of plausible criteria is inherently open-ended and subjective, False Positives are hard to define, so we use TPR as our primary metric: higher scores indicate more complete coverage of the reference annotated criteria.
Evaluation metrics for substructure uncovering. We measure how well each criterion-specific substructure predicted by the Grouper matches the ground truth along two complementary axes: i) Semantic consistency: We quantify the semantic agreement between predicted cluster names and ground-truth labels by computing their cosine similarity with Sentence-BERT (Reimers and Gurevych, 2019) for each image; the dataset-level mean is reported as Semantic Accuracy (SAcc) (Conti et al., 2023; Liu et al., 2024e). ii) Structural consistency: We quantify the structural agreement between predicted and ground-truth clusters with clustering accuracy (CAcc) (Vaze et al., 2022), which computes Hungarian-matched accuracy between the two partitions (Kuhn, 1955). Because ground-truth granularity is unknown during OpenSMC evaluation, we select the predicted substructure with the highest CAcc for evaluation. Unlike TCMC methods, which use the true cluster count during both training and evaluation, our protocol offers a fair and practical benchmark for OpenSMC.
Figure 4: Comprehensiveness Comparison of Criteria Proposers: TPR performance of each proposer is evaluated against Basic and Hard ground-truth criteria, and visualized using a Progress Bar Chart. Each block represents one ground-truth criterion, with Colored blocks indicating successfully discovered criteria and Gray blocks representing undiscovered criteria.
@@ 5.1 Study of the Criteria Proposer
Expanding ground-truth criteria for rigorous evaluation. The annotated basic criteria may not capture all valid groupings in complex datasets such as COCO-4c. We engaged human annotators to expand additional criteria for the six benchmarks in § 3, obtaining {10, 4, 11, 7, 17, 11} criteria for {Fruit-2c, Card-2c, Action-3c, Clevr-4c, COCO-4c, Food-4c}, respectively. We call the original per-image annotated criteria (Figure 2) the Basic ground truth and the enriched set the Hard ground truth. Full annotation details appear in App. C.2.
Which Proposer performs best? Figure 4 compares the TPR performance for all Proposer variants. The Caption-based Proposer recovers the most comprehensive set of criteria, ranking closest to the human-annotated reference set on every benchmark and on both Basic and Hard splits; in the latter it outperforms the runner-up Tag-based Proposer by $+32.2\%$ TPR. Captions offer richer, nuanced cues that help the LLM discover grouping criteria more comprehensively, whereas tags provide little context and therefore struggle on complex datasets (e.g., COCO-4c). The Image-based Proposer lags behind, as reasoning over small image grids omits fine visual cues and obscures overarching themes.
Impact of image quantity on criterion discovery. Figure 5 shows the TPR performance of the Caption-based Proposer as the number of input images for criteria discovery varies. Interestingly, for object-centric datasets (Card-2c, Clevr-4c), only a handful of samples – even a single image – recover most criteria, thanks to their uniform structure, i.e., seeing one playing card is enough to suggest criteria like Suit and Rank in Card-2c. In contrast, performance drops sharply on the more diverse and realistic datasets (COCO-4c, Food-4c, Action-3c) when fewer images are provided, because inferring their richer and intricate thematic criteria requires wider visual coverage. Since $\mathcal{X}$-Cluster assumes no dataset priors, we use all images by default.
Figure 5: Impact of Image Quantity on Criteria Discovery. We evaluate the TPR performance of the Caption-based Proposer at different image scales against the Hard ground-truth criteria set.
Figure 6: Comparison of Semantic Groupers. We report CAcc, SAcc, and their Harmonic Mean (HM) for different Semantic Groupers (Refer to caption) on the Basic criteria across six benchmarks. CLIP zero-shot classification (Refer to caption) serves as a pseudo upper bound, while KMeans (Refer to caption) with strong visual features is used as a CAcc baseline. Best performers are marked with Refer to caption. See App. G.2 for expanded numerical results and App. H for clustering visualizations.
@@
Table 2: Comparison with TCMC methods. For each benchmark, we report the average CAcc (%) and SAcc (%) across all criteria. We provide CLIP L/14 zero-shot performance as the pseudo upper-bound reference (UB). Note: -marked methods used the ground-truth criteria and the number of clusters ($K_{l}$) as prior input. MMaP and MSub do not build semantic clusters. See expanded results in App. G.3.
@@
SSD-LLM  (Luo et al., 2024b) 41.6 52.1 47.5 55.5 54.8 37.6 78.1 52.9 67.3 76.3 62.0 46.8 58.6 53.6
$\mathcal{X}$-Cluster (Ours) 51.2 48.4 48.1 64.9 64.9 54.3 68.3 60.6 73.3 84.3 65.1 61.1 61.8 62.3
5.2 Study of the Semantic Grouper
Which Semantic Grouper performs best? In Figure 6, we score each Grouper with CAcc (structural agreement) and SAcc (semantic agreement) evaluated on the ground-truth per-criterion annotation, ranking models by their harmonic mean Harmonic Mean (HM) of CAcc and SAcc. To gauge headroom we include a pseudo upper bound: CLIP ViT-L/14 in a zero-shot regime where the true criterion, cluster names and count are all supplied. We also report $k$-means baselines on features from CLIP-L/14, DINOv1-B/16 (Li et al., 2023b), and DINOv2-G/14 (Oquab et al., 2024), evaluated with CAcc and given the oracle $k$.
Figure 6 shows the Caption-based Grouper performs best. It ranks first on 12/19 criteria and averages $59.9\%$ CAcc, virtually matching the upper bound’s $58.1\%$. Its mean SAcc is $60.5\%$, again best among learned methods but below the upper-bound $74.2\%$, which benefits from exact ground-truth names. This gap is expected given lexical variability, e.g., “Joyful,” “Happy,” and “Cheerful” all describe the Mood criterion but lack full semantic equivalence. The BLIP-2 Image-based Grouper ranks second. Its criterion-specific questions help, yet per-image answers add noise. The Tag-based Grouper trails across all benchmarks, reflecting the weak alignment between generated tags and dataset concepts.
Effect of multi-granularity refinement. We ablate the Caption-based Grouper with three cluster naming schemes to study the effect of our multi-granularity design: i) Initial Names, where the initially LLM-assigned names are used as the final output; ii) Flat Refinement, where the LLM refines initial names into a single-level list with uniform granularity; and our iii) Multi-Granularity Refinement strategy. As Figure 7 indicates, either refinement method markedly boosts accuracy over raw initial names, confirming the need for consistent granularity. Furthermore, our multi-granularity strategy outperforms the flat one by offering several detail levels, providing greater flexibility in aligning with user-preferred grouping granularity.
Figure 7: Study of multi-granularity refinement.
5.3 Comparison with TCMC Methods
@@
Additional ablations and studies of $\bm{\mathcal{X}}$-Cluster. In the technical appendix, we provide extensive ablations and follow-up experiments: i) Cluster visualizations and qualitative analysis (App. H); ii) Failure case analysis (App. I); iii) Handling of hallucinated criteria (App. K); iv) Mitigating model biases (App. L); v) Computational cost analysis (App. M); vi) Sensitivity analysis of different MLLMs and LLMs (App. N); vii) Further investigation on multi-granularity clustering (App. J); and (viii) Fine-grained criterion handling (App. O).
Figure 8: Bias Discovery in T2I-Generated Images. Bias intensity, dominant clusters, and example images are shown for few occupations. Full results for all studied occupations are provided in App. P.1.